# IP-LLM: Semantic Space Informed Prompt Learning with LLM for Time Series Forecasting

### v1

@@ Abstract
Recently, there has been a growing interest in leveraging pre-trained large language models (LLMs) for various time series applications. However, the semantic space of LLMs, established through the pre-training, is still underexplored and may help yield more distinctive and informative representations to facilitate time series forecasting. To this end, we propose Semantic Space Informed Prompt learning with LLM (IP-LLM) to align the pre-trained semantic space with time series embeddings space and perform time series forecasting based on learned prompts from the joint space. We first design a tokenization module tailored for cross-modality alignment, which explicitly concatenates patches of decomposed time series components to create embeddings that effectively encode the temporal dynamics. Next, we leverage the pre-trained word token embeddings to derive semantic anchors and align selected anchors with time series embeddings by maximizing the cosine similarity in the joint space. This way, IP-LLM can retrieve relevant semantic anchors as prompts to provide strong indicators (context) for time series that exhibit different temporal dynamics. With thorough empirical studies on multiple benchmark datasets, we demonstrate that the proposed IP-LLM can achieve superior forecasting performance over state-of-the-art baselines. Furthermore, our ablation studies and visualizations verify the necessity of prompt learning informed by semantic space.
@@ 1 Introduction
Recently, significant efforts have been made to build foundation models for general-purpose time series analysis (Wu et al., 2023; Garza & Mergenthaler-Canseco, 2023; Rasul et al., 2023). TimesNet (Wu et al., 2023) uses TimesBlock as a task-general backbone that captures multi-periodicity and extracts complex intraperiod- and interperiod-variations via transformed 2D tensors. TimeGPT-1 describes a general pre-trained model for time series forecasting (Garza & Mergenthaler-Canseco, 2023). These approaches, however, are hindered by two main challenges. First, time series data can be acquired in various formats, such as univariate or multivariate, often in large volumes, and from different domains, like healthcare, finance, traffic, environmental sciences, etc. This escalates the complexity of model training and poses challenges in handling different scenarios. Second, time series data, in practice, often exhibit non-stationary characteristics, resulting in the underlying statistical properties, such as means, variances, and auto-correlations shifting during collection. This could also result in concept drift, where the statistical properties of target variables change over time. These realities present significant challenges for large models to be adapted and retrained effectively.
On the other hand, LLMs trained on extensive and diverse text corpora can serve as a foundational knowledge base that can be applied to a variety of downstream tasks with minimal task-specific prompt learning or fine-tuning. Inspired by this, there has been a growing interest in leveraging existing LLMs to facilitate time series analysis. For instance, Tian Zhou & Jin (2023) utilizes a frozen pre-trained language model to attain state-of-the-art or equivalent performance. Jin et al. (2024) develop time-LLM to reprogram the input time series via text prototype representations by incorporating the embeddings of the dataset’s text descriptions as context information. In real-world applications, however, dataset description information may not always be available or informative. In addition, the patching operation (i.e., tokenization), which splits a long time series sequence into overlapping segments over instance normalized time series input, may have limited expressibility as it could fail to capture the subtle variations in time series.
@@ 2.1 Time Series Forecasting
In recent years, a variety of statistical and machine learning methods have been developed for time series analysis, e.g., ARIMA (Anderson & Kendall, 1976), Prophet (Taylor & Letham, 2018), etc. More recently, different types of deep neural networks have been applied for time series analysis. For instance, recurrent neural network (RNN) based models have been developed to capture auto-regressive temporal dynamics (Li et al., 2017; Lai et al., 2018; Gu et al., 2021). Graph neural networks (GNN) based methods are leveraged to capture variable dependencies among different time series (Cao et al., 2020; Wu et al., 2020; Shang et al., 2021). Transformer based models leverage the self-attention mechanisms tailored for time series to better capture the temporal dynamics, variable dependencies, or both (Woo et al., 2022; Zhou et al., 2021; Wu et al., 2021; Zhou et al., 2022; Liu et al., 2023b). More recently, MLP-based models (Challu et al., 2023; Zeng et al., 2023) and convolution-based models (Wu et al., 2023) have achieved state-of-the-art performance on par with Transformers, but with much simpler designs. Nevertheless, while these deep forecasters perform well on specific datasets, they lack the flexibility and generalizability to a variety of real-world time series data.
@@ 2.2 Pre-trained Large Model for Time Series Analysis
Recent advancements in natural language processing (NLP) and computer vision (CV) demonstrate that pre-trained models can effectively adapt to a range of downstream tasks through fine-tuning (Bao et al., 2021; He et al., 2022; Brown et al., 2020; Devlin et al., 2018). Inspired by this, several different time series pre-trained models have been developed based on either supervised (Fawaz et al., 2018) or self-supervised learning (Zhang et al., 2022b; Deldari et al., 2022). During the training stage, models can learn robust representations from a variety of input time series data. Then, these models can be fine-tuned for downstream tasks of similar domains to enhance their performance (Tang et al., 2022). With the emergence and success of Large Language Models (LLMs), including T5 (Raffel et al., 2020), GPT-based models (Radford et al., 2018, 2019; Brown et al., 2020; Ouyang et al., 2022), and LLaMA (Touvron et al., 2023a), which have showcase their robust pattern recognition and reasoning abilities over complex sequences of tokens, there is a trend to explore how to effectively transfer knowledge from these powerful pre-trained LLM models to time series domain. One line of research focuses on leveraging the pre-trained LLMs as zero-shot learners. For instance, Xue & Salim (2022) and Nate Gruver & Wilson (2023) directly convert time series data to corresponding text sequence inputs and achieve encouraging results for time series forecasting. Another line of research (Tian Zhou & Jin, 2023; Chang et al., 2023) involves tokenizing the input time series data into overlapping patches and strategically leveraging or fine-tuning LLMs for time series analysis. Following this paradigm, TEST (Sun et al., 2023) and Time-LLM (Jin et al., 2024) reprogram time series data with text prototype embedding and incorporate textual prompts for time series analysis. TEMPO (Cao et al., 2023) incorporates the decomposition of time series and retrieval-based prompt design for non-stationary time series data. Different from those methods, we explicitly leverage semantic anchors derived from pre-trained word token embeddings (semantic space) to align time series embeddings and develop a simple yet effective prompt to inform LLM for forecasting tasks.
@@ 3 Methodology
Overview: IP-LLM consists of three key components as shown in Figure 2. Given the input time series, we first tokenize it and obtain the time series (TS) embedding based on Seasonal-Trend decomposition using Loess (STL) (Cleveland et al., 1990). Next, we will align the TS embedding with semantic anchors derived from the pre-trained word token embedding. Finally, top-k similar semantic anchors will be retrieved to serve as prefix-prompts for the TS embedding and the concatenated vector will be leveraged as the query for pre-trained LLMs.
In this paper, GPT-2 is used as the backbone. During the training stage, we not only learn the mapping functions of input and output but also fine-tune the positional embedding and layer norm block of GPT-2. During the test stage, all parameters are frozen to provide the forecasting results.
@@ 3.2 Time Series Tokenization
In many real-world applications, non-stationary data is prevalent. To tackle this problem, we adopt Seasonal-Trend decomposition using Loess (STL) (Cleveland et al., 1990), which is a robust method to decompose time series into long-term trend, seasonal, and residual components. Given the -th time series input at timestamp , i.e., , the additive STL decomposition is represented as: , where denotes the long-term trend, seasonal, and residual components, respectively. After the decomposition, we apply the reversible instance normalization (Kim et al., 2021) on each component such that each component has zero mean and unit standard deviation to mitigate the distribution shift in time series. Specifically, for the trend component of , i.e., , its transformed value can be given by:
where and are the instance-specific mean and variance, respectively. and are trainable parameters for this component. This transformation is also applied to both the seasonal and residual components. Next, we follow (Nie et al., 2023) to encode temporal information and local contexts of input time series by aggregating consecutive timestamps into overlapped patched tokens. Take the trend component as an example, the normalized component series, is converted to patched token representation , in which is the patch length, is the number of patches and is the horizontal sliding stride. We apply patching to each variable component over the temporal dimension and then concatenate the tokens of these components into a single meta-token, . We then feed the meta-token into a projection layer to get the time series embedding , where is the embedding size for the pre-trained LLMs.
@@ 3.3 Semantic Space Informed Prompting
Prompting has emerged as an effective technique in various applications, enabling LLMs to utilize task-specific information to achieve enhanced reasoning capabilities (Yin et al., 2023). Existing works primarily focus on employing template-based and fixed prompts for pre-trained LLMs in time series analysis (Xue & Salim, 2022; Jin et al., 2024). While these methods are intuitive, straightforward, and yield satisfactory results, their rigid prompt contexts are in line with linguistic semantics. However, time series representation inherently lacks human semantics and is more closely tied to sequence patterns in the form of temporal dynamics. Conversely, Lester et al. (2021) demonstrates the effectiveness of soft prompts in enabling LLMs to more effectively comprehend inputs. In the realm of time series analysis with LLMs, recent works (Sun et al., 2023; Cao et al., 2023) start to consider soft prompts as task-specific, randomly initialized, trainable vectors that learn from the supervised loss between LLM’s output and the ground truth. However, the semantic space of LLMs, established through the pre-training, is still underexplored and may help yield more distinctive and informative representations for time series data. Based on this intuition, we introduce a prompting mechanism informed by the pre-trained semantic space. Specifically, the pre-trained semantic word token embeddings, represented as where is the vocabulary size, are inevitably large and dense. For example, the vocabulary size of GPT-2 (Radford et al., 2019) reaches 50257 and may raise computational deficiency. Instead of directly using the semantic word token embedding, we derive a small set of semantic anchors in the hidden space using a generic mapping function on , which is denoted as , where is the reduced number of semantic anchors and . To properly retrieve relevant semantic anchors to enhance the time series embedding , we align the semantic anchors and time series embedding based on a score-matching function . In this paper, we implement the score-matching function based on cosine similarity:
@@
which will serve as the input for the pre-trained LLMs
@@ 3.4 Optimization Objective
We can obtain the output embedding after the forward path of the prompt enhanced time series embedding through LLMs. We will flatten it and use a linear mapping to project the representation to the forecasting horizon . The overall forecasting should also be the additive combination of the individual component predictions due to the decomposition step. We further split and express into a concatenation form and obtain the forecasting results as . At every training step, the overall training objective is:
where the first term is the forecasting loss in the form of mean squared error (MSE), and the second term is a score-matching function to align selected semantic anchors with the time series embedding obtained via STL decomposition and patching. In this way, we could obtain a more informative space to facilitate the underlying forecasting task. is a hyper-parameter to trade-off the alignment.
@@ 4 Experiments
In our experiments, we compare the proposed IP-LLM against a variety of baselines over 11 public datasets. We validate the effectiveness of IP-LLM over different time series tasks, including long-term forecasting (Section 4.1), short-term forecasting (Section 4.2), and few-shot forecasting (Section 4.3). We also provide the ablation studies and parameter sensitivity analysis in Section 4.4. Finally, we visualize the prompt enhanced time series embeddings to qualitatively assess the effectiveness of IP-LLM. We follow the experimental configurations (Wu et al., 2023) for all baselines using the unified pipeline111https://github.com/thuml/Time-Series-Library
@@
Baselines. The baselines include a set of Transformer-based methods, i.e., iTransformer (Liu et al., 2023b), PatchTST (Nie et al., 2023), FEDformer (Zhou et al., 2022), Autoformer (Wu et al., 2021), Non-Stationary Transformer (Liu et al., 2022), ETSformer (Woo et al., 2022) and Informer (Zhou et al., 2021). We also select a set of non-transformer based techniques, i.e., DLinear (Zeng et al., 2023), TimesNet (Wu et al., 2023), and LightTS (Zhang et al., 2022a) for comparison. Finally, two approaches based on LLMs, i.e., Time-LLM (Jin et al., 2024) and OFA (Tian Zhou & Jin, 2023) are also included. More detailed descriptions of baseline methods are provided in Appendix A.2

### v2

@@ Abstract
Recently, there has been a growing interest in leveraging pre-trained large language models (LLMs) for various time series applications. However, the semantic space of LLMs, established through the pre-training, is still underexplored and may help yield more distinctive and informative representations to facilitate time series forecasting. To this end, we propose Semantic Space Informed Prompt learning with LLM (IP-LLM) to align the pre-trained semantic space with time series embedding space and perform time series forecasting based on learned prompts from the joint space. We first design a tokenization module tailored for cross-modality alignment, which explicitly concatenates patches of decomposed time series components to create embeddings that effectively encode the temporal dynamics. Next, we leverage the pre-trained word token embeddings to derive semantic anchors and align selected anchors with time series embeddings by maximizing the cosine similarity in the joint space. This way, IP-LLM can retrieve relevant semantic anchors as prompts to provide strong indicators (context) for time series that exhibit different temporal dynamics. With thorough empirical studies on multiple benchmark datasets, we demonstrate that the proposed IP-LLM can achieve superior forecasting performance over state-of-the-art baselines. Furthermore, our ablation studies and visualizations verify the necessity of prompt learning informed by semantic space.
@@ 1 Introduction
Recently, significant efforts have been made to build foundation models for general-purpose time series analysis (Wu et al., 2023; Garza & Mergenthaler-Canseco, 2023; Rasul et al., 2023). TimesNet (Wu et al., 2023) uses TimesBlock as a task-general backbone to capture multi-periodicity and extract complex intraperiod- and interperiod-variations via transformed 2D tensors. TimeGPT-1 describes a general pre-trained model for time series forecasting (Garza & Mergenthaler-Canseco, 2023). These approaches, however, are hindered by two main challenges. First, time series data can be acquired in various formats, such as univariate or multivariate, often in large volumes, and from different domains, like healthcare, finance, traffic, environmental sciences, etc. This escalates the complexity of model training and poses challenges in handling different scenarios. Second, time series data, in practice, often exhibit non-stationary characteristics, resulting in the underlying statistical properties, such as means, variances, and auto-correlations shifting during collection. This could also result in concept drift, where the statistical properties of target variables change over time. These realities present significant challenges for large models to be adapted and retrained effectively.
On the other hand, LLMs trained on extensive and diverse text corpora can serve as a foundational knowledge base that can be applied to a variety of downstream tasks with minimal task-specific prompt learning or fine-tuning. Inspired by this, there has been a growing interest in leveraging existing LLMs to facilitate time series analysis. For instance, Tian Zhou & Jin (2023) utilizes a frozen pre-trained language model to attain state-of-the-art or equivalent performance. Jin et al. (2024) develop time-LLM to reprogram the input time series via text prototype representations by incorporating the embeddings of the dataset’s text descriptions as context information. In real-world applications, however, dataset description information may not always be available or informative. In addition, the patching operation (i.e., tokenization), which splits a long time series sequence into overlapping segments over instance normalized time series input, may have limited expressibility as it could fail to capture the subtle variations of different components in time series.
@@ 2.1 Time Series Forecasting
In recent years, a variety of statistical and machine learning methods have been developed for time series analysis, e.g., ARIMA (Anderson & Kendall, 1976), Prophet (Taylor & Letham, 2018), etc. More recently, different types of deep neural networks have been applied for time series analysis. For instance, recurrent neural network (RNN) based models have been developed to capture auto-regressive temporal dynamics (Qin et al., 2017; Li et al., 2017; Lai et al., 2018; Gu et al., 2021). Graph neural networks (GNN) based methods are leveraged to capture variable dependencies among different time series (Cao et al., 2020; Wu et al., 2020; Shang et al., 2021; Pan et al., 2024). Transformer based models leverage the self-attention mechanisms tailored for time series to better capture the temporal dynamics, variable dependencies, or both (Woo et al., 2022; Zhou et al., 2021; Wu et al., 2021; Zhou et al., 2022; Liu et al., 2023b). More recently, MLP-based models (Challu et al., 2023; Zeng et al., 2023) and convolution-based models (Wu et al., 2023) have achieved state-of-the-art performance on par with Transformers, but with much simpler designs. Nevertheless, while these deep forecasters perform well on specific datasets, they lack the flexibility and generalizability to adapt to real-world time series data from different domains.
@@ 2.2 Pre-trained Large Model for Time Series Analysis
Recent advancements in natural language processing (NLP) and computer vision (CV) demonstrate that pre-trained models can effectively adapt to a range of downstream tasks through fine-tuning (Bao et al., 2021; He et al., 2022; Brown et al., 2020; Devlin et al., 2018). Inspired by this, several different pre-trained models have been developed for time series based on either supervised (Fawaz et al., 2018) or self-supervised learning (Zhang et al., 2022b; Deldari et al., 2022). During the training stage, models can learn robust representations from a variety of input time series data. Then, these models can be fine-tuned for downstream tasks of similar domains to further enhance their performance (Tang et al., 2022). With the emergence and success of Large Language Models (LLMs), including T5 (Raffel et al., 2020), GPT-based models (Radford et al., 2018, 2019; Brown et al., 2020; Ouyang et al., 2022), and LLaMA (Touvron et al., 2023a), which have showcase their robust pattern recognition and reasoning abilities over complex sequences of tokens, there is a trend to explore how to effectively transfer knowledge from these powerful pre-trained LLM models to time series domain (Jiang et al., 2024). One line of research focuses on leveraging the pre-trained LLMs as zero-shot learners. For instance, Xue & Salim (2022) and Nate Gruver & Wilson (2023) directly convert time series data to corresponding text sequence inputs and achieve encouraging results for time series forecasting. Another line of research (Tian Zhou & Jin, 2023; Chang et al., 2023) involves tokenizing the input time series data into overlapping patches and strategically leveraging or fine-tuning LLMs for time series analysis. Following this paradigm, TEST (Sun et al., 2023) and Time-LLM (Jin et al., 2024) reprogram time series data with text prototype embedding and incorporate textual prompts for time series analysis. TEMPO (Cao et al., 2023) incorporates the decomposition of time series and retrieval-based prompt design for non-stationary time series data. Different from those methods, we explicitly leverage semantic anchors derived from pre-trained word token embeddings (semantic space) to align time series embeddings and develop a simple yet effective prompt mechanism to inform LLM for forecasting tasks.
@@ 3 Methodology
Overview: IP-LLM consists of three key components as shown in Figure 2. Given the input time series, we first tokenize it and obtain the time series (TS) embedding based on time series decomposition and patching. Next, we will align the TS embedding with semantic anchors derived from the pre-trained word token embedding. Finally, top- similar semantic anchors will be retrieved to serve as prefix-prompts for the TS embedding and the concatenated vector will be leveraged as the query for pre-trained LLMs.
In this paper, GPT-2 is used as the backbone. During the training stage, we not only learn the mapping functions of input and output but also fine-tune the positional embedding and layer norm block of GPT-2.
@@ 3.2 Time Series Tokenization
In real-world applications, non-stationary data is prevalent. To tackle this problem, we first apply the reversible instance normalization (Kim et al., 2021) on time series input such that the data has zero mean and unit standard deviation to mitigate the distribution shift in time series. Specifically, given the -th time series input at time step , i.e., , the transformed value can be given by:
where and are the instance-specific mean and variance, respectively. and are trainable parameters. Next, we adopt an additive seasonal-trend decomposition method to decompose normalized time series into long-term trend, seasonal, and residual components. The additive seasonal-trend decomposition is given by , where denotes the long-term trend, seasonal, and residual component, respectively. There are several options for additive seasonal-trend decomposition. One option is the classical additive seasonal-trend decomposition that first obtains long-term trend components using moving averages. Then, the seasonal component is estimated by averaging the detrended time series with pre-defined season parameters. Finally, the residual component is obtained by subtracting the estimated trend and seasonal components from the normalized time series. Another option is the Seasonal-Trend decomposition using Loess (STL) (Cleveland et al., 1990). The choice of decomposition method will based on validation results.
Next, we follow (Nie et al., 2023) to encode temporal information and local contexts of input time series by aggregating consecutive time steps into overlapped patched tokens. Take the trend component as an example, the normalized component series, is converted to patched token representation , in which is the patch length, is the number of patches and is the horizontal sliding stride. We apply patching to each variable component over the temporal dimension and then concatenate the tokens of these components into a single meta-token, . We then feed the meta-token into a projection layer to get the time series embedding , where is the embedding size for the pre-trained LLMs.
@@ 3.3 Semantic Space Informed Prompting
Prompting has emerged as an effective technique in various applications, enabling LLMs to utilize task-specific information to achieve enhanced reasoning capabilities (Yin et al., 2023). Existing works primarily focus on employing template-based and fixed prompts for pre-trained LLMs in time series analysis (Xue & Salim, 2022; Jin et al., 2024). While these methods are intuitive, straightforward, and yield satisfactory results, their rigid prompt contexts are in line with linguistic semantics. However, time series representation inherently lacks human semantics and is more closely tied to sequence patterns in the form of temporal dynamics. Conversely, Lester et al. (2021) demonstrate the effectiveness of soft prompts in enabling LLMs to comprehend inputs more effectively. In the realm of time series analysis with LLMs, recent works (Sun et al., 2023; Cao et al., 2023) start to consider soft prompts as task-specific, randomly initialized, trainable vectors that learn from the supervised loss between LLM’s output and the ground truth. However, the semantic space of LLMs, established through the pre-training, is still underexplored and may help yield more distinctive and informative representations for time series data. Based on this intuition, we introduce a prompting mechanism informed by the pre-trained semantic space. Specifically, the pre-trained semantic word token embeddings, represented as where is the vocabulary size, are inevitably large and dense. For example, the vocabulary size of GPT-2 (Radford et al., 2019) reaches 50,257 and may raise computational deficiency. Instead of directly using the semantic word token embedding, we derive a small set of semantic anchors in the hidden space using a generic mapping function on , which is denoted as , where is the reduced number of semantic anchors and . To properly retrieve relevant semantic anchors to enhance the time series embedding , we align the semantic anchors and time series embedding based on a score-matching function . In this paper, we implement the score-matching function based on cosine similarity
@@
which will serve as the input for the pre-trained LLMs.
@@ 3.4 Optimization Objective
We can obtain the output embedding after the forward path of the prompt enhanced time series embedding through LLMs. We will flatten it and use a linear mapping to project the representation to the forecasting horizon . The overall forecasting should also be the additive combination of the individual component predictions due to the decomposition step. We further split and express into a concatenation form and obtain the forecasting results as . At every training iteration, the overall training objective is:
where the first term is the forecasting loss in the form of mean squared error (MSE), and the second term is a score-matching function to align selected semantic anchors with the time series embedding obtained via decomposition and patching. In this way, we could obtain a more informative space to facilitate the underlying forecasting task. is a hyper-parameter to trade-off the alignment.
@@ 4 Experiments
In our experiments, we compare the proposed IP-LLM against a variety of baselines on 11 public datasets. We validate the effectiveness of IP-LLM over different time series tasks, including long-term forecasting (Section 4.1), short-term forecasting (Section 4.2), and few-shot forecasting (Section 4.3). We also provide the ablation studies and parameter sensitivity analysis in Section 4.4. Finally, we visualize the prompt enhanced time series embeddings to qualitatively assess the effectiveness of IP-LLM. We follow the experimental configurations (Wu et al., 2023) for all baselines using the unified pipeline.111https://github.com/thuml/Time-Series-Library
@@
Baselines. The baselines include a set of Transformer-based methods, i.e., iTransformer (Liu et al., 2023b), PatchTST (Nie et al., 2023), FEDformer (Zhou et al., 2022), Autoformer (Wu et al., 2021), Non-Stationary Transformer (Liu et al., 2022), ETSformer (Woo et al., 2022) and Informer (Zhou et al., 2021). We also select a set of non-transformer based techniques, i.e., DLinear (Zeng et al., 2023), TimesNet (Wu et al., 2023), and LightTS (Zhang et al., 2022a) for comparison. Finally, two approaches based on LLMs, i.e., OFA (Tian Zhou & Jin, 2023) and Time-LLM (Jin et al., 2024)222We reproduced results through public available code: https://github.com/KimMeen/Time-LLM. For Time-LLM variants, ‘L’ denotes the LLaMA backbone, and ‘G’ refers to the GPT-2 backbone..
| Methods | IP-LLM | Time-LLM(L) | Time-LLM(G) | OFA | iTransformer | Dlinear | PatchTST | ||||||||
| Weather | 96 | 0.145 | 0.195 | 0.148 | 0.197 | 0.158 | 0.210 | 0.162 | 0.212 | 0.253 | 0.304 | 0.176 | 0.237 | 0.149 | 0.198 |
| 192 | 0.190 | 0.235 | 0.194 | 0.246 | 0.197 | 0.245 | 0.204 | 0.248 | 0.280 | 0.319 | 0.220 | 0.282 | 0.194 | 0.241 | |
| 336 | 0.243 | 0.280 | 0.248 | 0.285 | 0.248 | 0.285 | 0.254 | 0.286 | 0.321 | 0.344 | 0.265 | 0.319 | 0.245 | 0.282 | |
| 720 | 0.312 | 0.326 | 0.317 | 0.332 | 0.319 | 0.334 | 0.326 | 0.337 | 0.364 | 0.374 | 0.333 | 0.362 | 0.314 | 0.334 | |
| Avg. | 0.222 | 0.259 | 0.226 | 0.265 | 0.230 | 0.268 | 0.237 | 0.270 | 0.304 | 0.335 | 0.248 | 0.300 | 0.225 | 0.264 | |
| Electricity | 96 | 0.135 | 0.230 | 0.140 | 0.246 | 0.137 | 0.237 | 0.139 | 0.238 | 0.147 | 0.248 | 0.140 | 0.237 | 0.129 | 0.222 |
| 192 | 0.149 | 0.247 | 0.155 | 0.253 | 0.150 | 0.249 | 0.153 | 0.251 | 0.165 | 0.267 | 0.153 | 0.249 | 0.157 | 0.240 | |
| 336 | 0.167 | 0.266 | 0.175 | 0.279 | 0.168 | 0.266 | 0.169 | 0.266 | 0.178 | 0.279 | 0.169 | 0.267 | 0.163 | 0.259 | |
| 720 | 0.200 | 0.287 | 0.204 | 0.305 | 0.203 | 0.293 | 0.206 | 0.297 | 0.322 | 0.398 | 0.203 | 0.301 | 0.197 | 0.290 | |
| Avg. | 0.161 | 0.257 | 0.168 | 0.270 | 0.164 | 0.261 | 0.167 | 0.263 | 0.203 | 0.298 | 0.166 | 0.263 | 0.161 | 0.252 | |
| Traffic | 96 | 0.379 | 0.274 | 0.383 | 0.280 | 0.380 | 0.277 | 0.388 | 0.282 | 0.367 | 0.288 | 0.410 | 0.282 | 0.360 | 0.249 |
| 192 | 0.397 | 0.282 | 0.399 | 0.294 | 0.399 | 0.288 | 0.407 | 0.290 | 0.378 | 0.293 | 0.423 | 0.287 | 0.379 | 0.256 | |
| 336 | 0.407 | 0.289 | 0.411 | 0.306 | 0.408 | 0.290 | 0.412 | 0.294 | 0.389 | 0.294 | 0.436 | 0.296 | 0.392 | 0.264 | |
| 720 | 0.440 | 0.301 | 0.448 | 0.319 | 0.445 | 0.308 | 0.450 | 0.312 | 0.401 | 0.304 | 0.466 | 0.315 | 0.432 | 0.286 | |
| Avg. | 0.405 | 0.286 | 0.440 | 0.301 | 0.408 | 0.290 | 0.414 | 0.294 | 0.389 | 0.295 | 0.433 | 0.295 | 0.390 | 0.263 | |
| ETTh1 | 96 | 0.366 | 0.396 | 0.380 | 0.406 | 0.383 | 0.410 | 0.379 | 0.402 | 0.395 | 0.420 | 0.367 | 0.396 | 0.379 | 0.407 |
| 192 | 0.401 | 0.420 | 0.426 | 0.438 | 0.419 | 0.435 | 0.415 | 0.424 | 0.427 | 0.441 | 0.401 | 0.419 | 0.428 | 0.442 | |
| 336 | 0.412 | 0.431 | 0.437 | 0.451 | 0.426 | 0.440 | 0.435 | 0.440 | 0.445 | 0.457 | 0.434 | 0.449 | 0.465 | 0.465 | |
| 720 | 0.440 | 0.458 | 0.515 | 0.509 | 0.428 | 0.456 | 0.441 | 0.459 | 0.537 | 0.530 | 0.472 | 0.493 | 0.504 | 0.500 | |
| Avg. | 0.406 | 0.427 | 0.439 | 0.451 | 0.414 | 0.435 | 0.418 | 0.431 | 0.451 | 0.462 | 0.418 | 0.439 | 0.444 | 0.453 | |
| ETTh2 | 96 | 0.278 | 0.340 | 0.306 | 0.362 | 0.297 | 0.357 | 0.289 | 0.347 | 0.304 | 0.360 | 0.301 | 0.367 | 0.296 | 0.353 |
| 192 | 0.346 | 0.385 | 0.346 | 0.385 | 0.349 | 0.390 | 0.358 | 0.392 | 0.377 | 0.403 | 0.394 | 0.427 | 0.382 | 0.404 | |
| 336 | 0.367 | 0.406 | 0.393 | 0.422 | 0.373 | 0.408 | 0.383 | 0.414 | 0.405 | 0.429 | 0.506 | 0.495 | 0.402 | 0.425 | |
| 720 | 0.400 | 0.436 | 0.397 | 0.433 | 0.400 | 0.436 | 0.438 | 0.456 | 0.443 | 0.464 | 0.805 | 0.635 | 0.444 | 0.465 | |
| Avg. | 0.347 | 0.391 | 0.360 | 0.400 | 0.355 | 0.398 | 0.367 | 0.402 | 0.382 | 0.414 | 0.502 | 0.481 | 0.381 | 0.411 | |
| ETTm1 | 96 | 0.288 | 0.346 | 0.311 | 0.365 | 0.291 | 0.346 | 0.296 | 0.353 | 0.312 | 0.366 | 0.304 | 0.348 | 0.303 | 0.351 |
| 192 | 0.323 | 0.365 | 0.364 | 0.395 | 0.336 | 0.373 | 0.335 | 0.373 | 0.347 | 0.385 | 0.336 | 0.367 | 0.341 | 0.376 | |
| 336 | 0.359 | 0.390 | 0.369 | 0.398 | 0.362 | 0.390 | 0.369 | 0.394 | 0.379 | 0.404 | 0.368 | 0.387 | 0.377 | 0.401 | |
| 720 | 0.403 | 0.418 | 0.416 | 0.425 | 0.410 | 0.421 | 0.418 | 0.424 | 0.441 | 0.442 | 0.421 | 0.418 | 0.431 | 0.436 | |
| Avg. | 0.343 | 0.379 | 0.365 | 0.395 | 0.349 | 0.382 | 0.355 | 0.386 | 0.370 | 0.399 | 0.357 | 0.389 | 0.363 | 0.391 | |
| ETTm2 | 96 | 0.165 | 0.257 | 0.170 | 0.262 | 0.184 | 0.275 | 0.170 | 0.264 | 0.179 | 0.271 | 0.168 | 0.263 | 0.173 | 0.262 |
| 192 | 0.222 | 0.299 | 0.229 | 0.303 | 0.238 | 0.310 | 0.231 | 0.306 | 0.242 | 0.313 | 0.229 | 0.310 | 0.231 | 0.300 | |
| 336 | 0.277 | 0.330 | 0.281 | 0.335 | 0.286 | 0.340 | 0.280 | 0.339 | 0.288 | 0.344 | 0.289 | 0.352 | 0.292 | 0.345 | |
| 720 | 0.363 | 0.390 | 0.379 | 0.403 | 0.379 | 0.403 | 0.373 | 0.402 | 0.378 | 0.397 | 0.416 | 0.437 | 0.371 | 0.394 | |
| Avg. | 0.257 | 0.319 | 0.264 | 0.325 | 0.271 | 0.332 | 0.265 | 0.328 | 0.272 | 0.331 | 0.275 | 0.340 | 0.267 | 0.325 | |